# Characterizing Dynamical Stability of Stochastic Gradient Descent in Overparameterized Learning

### v1

@@ 1 Introduction
Since the success of “AlexNet” [KSH12] in 2012, the overparameterization paradigm has become ubiquitous in modern machine learning. While it is well known that artificial neural networks with sufficiently many parameters can approximate arbitrary goal functions (see e.g. [HSW89]), the training process, especially for overparameterized networks, is not well understood and leaves many open questions. For one, the fact that the loss landscapes are usually non-convex, makes it difficult to rigorously guarantee that the optimization algorithms converge to global minima.
@@
Another key open problem in understanding the training of overparameterized networks is that of generalization (see e.g. [ZBH+21]). Classical wisdom suggests that overparameterized networks should suffer from overfitting problems due to their high expressivity. In practice, however, this is not the case. Put shortly, sufficiently large (i.e. deep, wide or both) artificial neural networks are capable of fitting arbitrary data, due to their large number of parameters (typically weights and biases). Thus, finding a set of parameters for which the network interpolates the training data does not in itself indicate that the network makes reasonable predictions for inputs outside the training data set. However, the parameters found by common optimization algorithms tend to perform well on unseen data. This is essential for the success of modern machine learning.
@@
One possible explanation for this phenomenon is that of dynamical stability (see e.g [WME18]). Since the learning rate is in practice not infinitesimally small, some global minima can become dynamically unstable, in the sense that the optimization algorithm will not converge to these solutions, even if it is initialized arbitrarily close to them. This excludes unstable optimal solutions from being found by the optimization algorithm, thereby reducing the effective hypothesis class to stable global minima.
@@
For gradient descent, a second order Taylor expansions shows that a global minimum is only stable if the largest eigenvalue of the Hessian of the empirical loss is less than $2/\eta$, where $\eta$ is the learning rate. Thus gradient descent can only converge to minima which are sufficiently flat, which has been been associated to good generalization (see e.g. [HS97]). Moreover, numerical experiments [CKL+20] have demonstrated that (deterministic) gradient descent operates at the edge of stability, meaning that the largest eigenvalue of the Hessian oscillates around the critical value $2/\eta$ for the later stages of training.
@@
Transferring these ideas to the stochastic case requires an appropriate notion of dynamical stability for stochastic gradient descent [WME18, MY21]. The goal of this work is to establish a coherent mathematical framework in which the dynamical stability of deterministic and stochastic optimization algorithms can be investigated and the convergence properties of global minima can be characterized.
@@ 1.1 Contributions
  • For global minima $x^{*}$, we introduce a new quantity, denoted by $\lambda(x^{*})$ (cf. (2.12) below), which can be seen as a characteristic Lyapunov exponent ([Ose68]) and corresponds to a new notion of dynamic stability/instability for SGD. It should be noted that this notion differs [^1] In particular our notion of linar stability is strictly weaker than the one introduced in [WME18, Defintion 1] (cf. Appendix A). from the ones introduced in [WME18] and [MY21] (cf. Appendix A).
@@
This clarifies the concept of dynamical stability for stochastic gradient descent and establishes a connection to a substantial body of mathematical and physical literature on Lyapunov exponents (see e.g. [Ose68, BL85, Arn98]). Moreover, it provides a foundation for new theoretic and numerical investigations of the learning process under stochastic gradient descent. Possible questions of interest include how the value of $\lambda$ evolves during training with stochastic gradient descent and which implications a small, respectively large, value of $\lambda$ has on generalization.
From a mathematical perspective our analysis of SGD amounts to a study of the asymptotic behavior of discrete-time random dynamical systems (see e.g. [Arn98] and also section 3.5) possessing a manifold of equilibria. In the unstable case, we use a method based on moment Lyapunov exponents [AKO86]. A more detailed discussion of this approach is given in section 3.1 an throughout the proofs.
@@
1.2 Literature
Closest to our work is a series of two papers [WME18, MY21], which studies the dynamical stability of stochastic gradient descent. We explain the relation to this work in detail in Appendix A. Convergence rates for SGD have been studied in
The most critical part of our proof based on techniques originally developed to study synchronization in stochastic differential equations [BS88, Bax91], which has seen recent attention motivated by fluid dynamics [CZH21, BBPS22, BCZG23].
The edge of stability phenomenon for gradient descent mentioned in the introduction has been numerically observed in [WME18, CKL+20] and theoretically studied in [ALP22]. Moreover, the difference between global minima found by GD and SGD [KMN+17, WME18] and SGD and Adam [KS17] has been investigated numerically.
The dynamics of stochastic gradient descent has also been studied using approximations by stochastic differential equations, called stochastic modified equations (see e.g. [LTE17, LTE19]). We note that such an approach is not compatible with our analysis, as it assumes an infinitesimally small learning rate. Moreover, stochastic modified equations have been extended to a stochastic flow in [GKK24], building a framework analogous to the one we introduce in section 3.5.
@@ 1.3 Organization of the Paper
We start by introducing the fundamental setting of the optimization task (section 2.1), as well as the learning algorithms we consider (2.2). In section 2.3, we heuristically derive the notions of linear stability for GD and SGD. These derivations are made rigorous in our main results, which we present in section 2.4. In section 2.5, the main quantities of interest, $\mu$ and $\lambda$, are related to the neural tangent kernel [JGH18]. Finally, we discuss possible generalizations of our main results in section 2.6. Section 3 provides the proofs of the main theorems. An overview over the structure of the proofs is given in section 3.1.
@@ 2.1 Network Model
As a consequence of Sard’s theorem (see e.g. [Mil97]), this condition will be satisfied for generic training data [^3] Put more precisely, for every smooth network model $\mathfrak{F}$ and every choice of $(y_{1},\dots,y_{N})$, there exists a set $\mathcal{Z}\subseteq\mathbb{R}^{N}$ of full Lebesgue measure, such that, if $(z_{1},\dots,z_{N})\in\mathcal{Z}$, then $\mathcal{M}$ is an embedded smooth $(D-N)$-dimensional manifold (cf. [Coo21]). and in the following we will assume that $\mathcal{M}$ is a smooth manifold as described (cf. Hypothesis (H1) below).
@@ 2.2 Learning Algorithms
In practice, a wide range of optimization algorithms (cf. e.g. [SSH21]) are used to find a global minimum $x^{*}\in\mathcal{M}$. In this work, we will restrict our study to the two most basic optimization algorithms: gradient descent (GD) and stochastic gradient descent (SGD). Both algorithms are initialized at a random point $X_{0}^{\operatorname{GD/SGD}}\in\mathbb{R}^{D}$, which is distributed according to some probability measure $\nu:\mathfrak{B}(\mathbb{R}^{D})\to[0,1]$, called initial distribution. [^4] Here and in the following, $\mathfrak{B}(\mathbb{R}^{D})$ denotes the family of Borel sets. We will assume that $\nu$ is equivalent to Lebesgue measure [^5] This is for example consistent with $\nu$ being a normal distribution, which is the most common choice in practice. (cf. Hypothesis (H2) below).
@@ 2.3 Linear Stability
where the last equivalence holds if $\eta<2$. While this formulation is more common in the literature (cf. e.g. [WME18, CKL+20, ALP22]), we will stick to the expression $\mu(x^{*})<0$ here, as it can be more easily extended to stochastic gradient descent.
@@
At first glance, this definition raises two problems. Firstly, it is not clear whether the limit exists. Secondly, the value of $\lambda(x^{*})$ seems to depend on the random sequence $(\xi_{1},\xi_{2},\dots)$. However, Kingman’s sub-additive ergodic theorem ([Kin68], see also [Ste89]) [^6] The integrability condition for Kingman’s ergodic theorem holds trivially here, as we only consider finitely man different matrices. states that for almost every realization $(\xi_{1},\xi_{2},\dots)$ we have
@@
We will call $x^{*}$ linearly stable (or linearly unstable) under stochastic gradient descent if $\lambda(x^{*})<0$ (or $\lambda(x^{*})>0$, respectively). [^7] As mentioned in the introduction, this definition differs from the ones introduced in [WME18] and [MY21]. See Appendix A for a detailed comparison.
@@ 2.4 Main Result
Recall that, by an argument in [Coo21], the Hypothesis (H1) is fulfilled for generic training data, while (H2) holds for the most common choice of initial distribution.
@@
Remark 2.2.
Both Theorem A and Theorem B fail to address the case $\mu(x^{*})=0$, respectively $\lambda(x^{*})=0$. Since the support of $X_{\lim}^{GD/SGD}$ is closed by definition and $\mu:\mathcal{M}\to\mathbb{R}$ is continuous, most points with $\mu(x^{*})=0$ should have global minima $x^{\prime}\in\mathcal{M}$ with $\mu(x^{\prime})<0$ nearby and thus be in the support of $X_{\lim}^{GD}$. Making this argument rigorous would require more precise knowledge of the network function $\mathfrak{F}$ and is outside the scope of this paper. Continuity of $\lambda$ is a more subtle issue. In general, $\lambda:\mathcal{M}\to\mathbb{R}$ is only upper semi-continuous [^8] This can readily be seen from the fact that it is defined as an infimum of continuous function.. However, a recent result by Avila, Esken and Viana [AEV23] shows that $\lambda$ is continuous at all points $x^{*}$, where the matrices $\mathcal{G}^{\prime}_{\eta,1}(x^{*}),\dots,\mathcal{G}^{\prime}_{\eta,N}(x^{*})$ are all invertible. This is, in particular, the case for all regular $x^{*}\in\mathcal{M}$.
@@ 2.5 Relation to the Neural Tangent Kernel
We point out that, for any $x^{*}\in\mathcal{M}$, the values of $\mu(x^{*})$ and $\lambda(x^{*})$ can be deduced entirely from the neural tangent kernel $K_{x^{*}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, first introduced in [JGH18], which is defined by
@@
This observation is useful for two reasons. On the one hand it connects the central quantities in the present work to an ongoing line of research of neural networks in the infinite width limit [JGH18]. On the other hand, it turns out that the expressions (2.18) and (2.19) are nicer to work with than the corresponding expressions in the previous section and will be exclusively used in the proofs of the main theorems.
@@ 2.6 Possible Extensions and Outlook
where $G_{x^{*},[\Xi]}$ denotes the Gram matrix $G_{x^{*}}$ of the neural tangent kernel with all rows whose index is not in $\Xi$ set to 0. In order to derive a version of Theorem B it is necessary to adapt the notion of regular global minima (cf. Definition 2.1) to ensure that the analog of Lemma 3.9 in the proof below still remains valid. We leave this as a problem for future work. The rest of the proof of Theorem B can be extended to mini-batch SGD with only minor changes. Such an extension of Theorem B could help to better understand the impact of the learning rate and the mini-batch size on generalization properties (see e.g. [HHS17, GDG+17, KMN+17] for numerical studies).
Furthermore it would be interesting to extend our analysis to other versions of stochastic gradient descent such as SGD with momentum [RHW86] or Adam [KB14]. The derivations in section 2.3 can be extended to these algorithms by linearizing the iteration steps around fixed points where all moment terms are zero. In the case of Adam, the parameter $\epsilon$ appearing in the numerator of the final update step is usually chosen extremely small ([KB14] suggests $\epsilon=10^{-8}$). Thus the linearization only approximates the dynamics of the actual algorithm in a vanishingly small neighborhood and it is questionable whether it is still meaningful for the dynamics of the algorithm.
@@
can, even for large $n$, still be negative with small, but positive probability. This effect is captured by a large deviation principle ([AK87]), the rate function of which depends on the moment Lyapunov exponents (see Appendix A for a definition).

### v2

@@ Abstract
Keywords: machine learning, stochastic gradient descent, overparametrization, linear stability, Lyapunov exponents
@@ 1 Introduction
Since the success of “AlexNet” (Krizhevsky et al. (2012)), the overparameterization paradigm has become ubiquitous in modern machine learning. While it is well known that artificial neural networks with sufficiently many parameters can approximate arbitrary goal functions (see e.g. Hornik et al. (1989)), the training process, especially for overparameterized networks, is not well understood and leaves many open questions. For one, the fact that the loss landscapes are usually non-convex, makes it difficult to rigorously guarantee that the optimization algorithms converge to global minima.
@@
Another key open problem in understanding the training of overparameterized networks is that of generalization (see e.g. Zhang et al. (2021)). Classical wisdom suggests that overparameterized networks should suffer from overfitting problems due to their high expressivity. In practice, however, this is not the case. Put shortly, sufficiently large (i.e. deep, wide or both) artificial neural networks are capable of fitting arbitrary data, due to their large number of parameters (typically weights and biases). Thus, finding a set of parameters for which the network interpolates the training data does not in itself indicate that the network makes reasonable predictions for inputs outside the training data set. However, the parameters found by common optimization algorithms tend to perform well on unseen data. This is essential for the success of modern machine learning.
@@
One possible explanation for this phenomenon is that of dynamical stability (see e.g Wu et al. (2018)). Since the learning rate is in practice not infinitesimally small, some global minima can become dynamically unstable, in the sense that the optimization algorithm will not converge to these solutions, even if it is initialized arbitrarily close to them. This excludes unstable optimal solutions from being found by the optimization algorithm, thereby reducing the effective hypothesis class to stable global minima.
@@
For gradient descent, a second order Taylor expansions shows that a global minimum is only stable if the largest eigenvalue of the Hessian of the empirical loss is less than $2/\eta$, where $\eta$ is the learning rate. Thus gradient descent can only converge to minima which are sufficiently flat, which has been been associated to good generalization (see e.g. Hochreiter and Schmidhuber (1997)). Moreover, numerical experiments by Cohen et al. (2020) have demonstrated that (deterministic) gradient descent operates at the edge of stability, meaning that the largest eigenvalue of the Hessian oscillates around the critical value $2/\eta$ for the later stages of training.
@@
Transferring these ideas to the stochastic case requires an appropriate notion of dynamical stability for stochastic gradient descent (cf. Wu et al. (2018) and Ma and Ying (2021)). The goal of this work is to establish a coherent mathematical framework in which the dynamical stability of deterministic and stochastic optimization algorithms can be investigated and the convergence properties of global minima can be characterized.
@@ 1.1 Contributions
  • For global minima $x^{*}$, we introduce a new quantity, denoted by $\lambda(x^{*})$ (cf. (2.12) below), which can be seen as a characteristic Lyapunov exponent (Oseledec (1968)) and corresponds to a new notion of dynamic stability/instability for SGD. It should be noted that this notion differs [^1] In particular our notion of linar stability is strictly weaker than the one introduced in (Wu et al., 2018, Defintion 1) (cf. Appendix A). from the ones introduced in Wu et al. (2018) and Ma and Ying (2021) (cf. Appendix A).
@@
This clarifies the concept of dynamical stability for stochastic gradient descent and establishes a connection to a substantial body of mathematical and physical literature on Lyapunov exponents (see e.g. Oseledec (1968); Bougerol and Lacroix (1985); Arnold (1998)). Moreover, it provides a foundation for new theoretic and numerical investigations of the learning process under stochastic gradient descent. Possible questions of interest include how the value of $\lambda$ evolves during training with stochastic gradient descent and which implications a small, respectively large, value of $\lambda$ has on generalization.
From a mathematical perspective our analysis of SGD amounts to a study of the asymptotic behavior of discrete-time random dynamical systems (see e.g. Arnold (1998) and also section 3.5) possessing a manifold of equilibria. In the unstable case, we use a method based on moment Lyapunov exponents (see e.g. Arnold et al. (1986)). A more detailed discussion of this approach is given in section 3.1 an throughout the proofs.
@@
1.2 Related Literature
Closest to our work is a series of two papers by Wu et al. (2018) and Ma and Ying (2021), which study the dynamical stability of stochastic gradient descent. We explain the relation to this work in detail in Appendix A.
Gurbuzbalaban et al. (2021) consider Lyapunov exponents of SGD in a setting where SGD does not converge to a global minimum, but the transition densities approach a heavy-tailed stationary distribution (see also e.g. Hodgkinson and Mahoney (2021)). Convergence rates for SGD with decaying learning rate have been studied in Fehrman et al. (2020).
The edge of stability phenomenon for gradient descent mentioned in the introduction has been numerically observed by Wu et al. (2018), as well as Cohen et al. (2020) and theoretically studied by Arora et al. (2022). Moreover, the difference between global minima found by GD and SGD (see Keskar et al. (2017); Wu et al. (2018)) and SGD and Adam (see Keskar and Socher (2017)) has been investigated numerically.
The dynamics of stochastic gradient descent has also been studied using approximations by stochastic differential equations, called stochastic modified equations (see e.g. Li et al. (2017, 2019)). We note that such an approach is not compatible with our analysis, as it assumes an infinitesimally small learning rate. Moreover, stochastic modified equations have been extended to a stochastic flow in Gess et al. (2024), building a framework analogous to the one we introduce in section 3.5.
The most critical part of our proof is based on techniques originally developed to study synchronization in stochastic differential equations by Baxendale and Stroock (1988) (see also Baxendale (1991)), which has seen recent attention motivated by fluid dynamics (Coti Zelati and Hairer (2021); Bedrossian et al. (2022); Blumenthal et al. (2023)).
@@ 1.3 Organization of the Paper
We start by introducing the fundamental setting of the optimization task (section 2.1), as well as the learning algorithms we consider (2.2). In section 2.3, we heuristically derive the notions of linear stability for GD and SGD. These derivations are made rigorous in our main results, which we present in section 2.4. In section 2.5, the main quantities of interest, $\mu$ and $\lambda$, are related to the neural tangent kernel (cf. Jacot et al. (2018)). Finally, we discuss possible generalizations of our main results in section 2.6. Section 3 provides the proofs of the main theorems. An overview over the structure of the proofs is given in section 3.1.
@@ 2.1 Network Model
As a consequence of Sard’s theorem (see e.g. Milnor (1997)), this condition will be satisfied for generic training data [^3] Put more precisely, for every smooth network model $\mathfrak{F}$ and every choice of $(y_{1},\dots,y_{N})$, there exists a set $\mathcal{Z}\subseteq\mathbb{R}^{N}$ of full Lebesgue measure, such that, if $(z_{1},\dots,z_{N})\in\mathcal{Z}$, then $\mathcal{M}$ is an embedded smooth $(D-N)$-dimensional manifold (cf. Cooper (2021)). and in the following we will assume that $\mathcal{M}$ is a smooth manifold as described (cf. Hypothesis (H1) below).
@@ 2.2 Learning Algorithms
In practice, a wide range of optimization algorithms (cf. e.g. Schmidt et al. (2021)) are used to find a global minimum $x^{*}\in\mathcal{M}$. In this work, we will restrict our study to the two most basic optimization algorithms: gradient descent (GD) and stochastic gradient descent (SGD). Both algorithms are initialized at a random point $X_{0}^{\operatorname{GD/SGD}}\in\mathbb{R}^{D}$, which is distributed according to some probability measure $\nu:\mathfrak{B}(\mathbb{R}^{D})\to[0,1]$, called initial distribution. [^4] Here and in the following, $\mathfrak{B}(\mathbb{R}^{D})$ denotes the family of Borel sets. We will assume that $\nu$ is equivalent to Lebesgue measure [^5] This is for example consistent with $\nu$ being a normal distribution, which is the most common choice in practice. (cf. Hypothesis (H2) below).
@@ 2.3 Linear Stability
where the last equivalence holds if $\eta<2$. While this formulation is more common in the literature (cf. e.g. Wu et al. (2018); Cohen et al. (2020); Arora et al. (2022)), we will stick to the expression $\mu(x^{*})<0$ here, as it can be more easily extended to stochastic gradient descent.
@@
At first glance, this definition raises two problems. Firstly, it is not clear whether the limit exists. Secondly, the value of $\lambda(x^{*})$ seems to depend on the random sequence $(\xi_{1},\xi_{2},\dots)$. However, Kingman’s sub-additive ergodic theorem (Kingman (1968), see also Steele (1989)) [^6] The integrability condition for Kingman’s ergodic theorem holds trivially here, as we only consider finitely many different matrices. states that for almost every realization $(\xi_{1},\xi_{2},\dots)$ we have
@@
In fact, the following theorem due to Oseledec (1968) (see also e.g. chapter 3.4 in Arnold (1998)) shows the existence of so-called Oseledec’s subspaces which are the equivalent notion to generalized eigenspaces for random matrix products.
Theorem 2.1 (Multiplicative Ergodic Theorem)
There exists a natural number $1\leq k\leq N$ and a tuple of extended real numbers
$-\infty\leq\lambda_{k}<\dots<\lambda_{2}<\lambda_{1}=\lambda(x^{*})$
and a tuple of multiplicities $m_{1},m_{2},\dots,m_{k}$ with $m_{1}+\dots+m_{k}=N$ such that the following holds: For almost every777with respect to the uniform i.i.d. measure, sequence $(\xi_{1},\xi_{2},\dots)$ there exists a tuple of subspaces
$\{0\}=V_{k+1}(\xi_{1},\xi_{2},\dots)\subsetneq V_{k}(\xi_{1},\xi_{2},\dots)\subsetneq\dots\subsetneq V_{2}(\xi_{1},\xi_{2},\dots)\subsetneq V_{1}(\xi_{1},\xi_{2},\dots)=\mathcal{N}(x^{*})$
with $\operatorname{dim}(V_{i})=m_{k}+\dots+m_{i}$ such that for every $1\leq i\leq k$
$\lim_{n\to\infty}\frac{1}{n}\log\left[\big{\|}\mathcal{G}^{\prime}_{\eta,\xi_{n}}(x^{*})\dots\mathcal{G}^{\prime}_{\eta,\xi_{1}}(x^{*})y\big{\|}\right]=\lambda_{i},~{}\forall\,y\in V_{i}(\xi_{1},\xi_{2},\dots)\setminus V_{i+1}(\xi_{1},\xi_{2},\dots).$
In particular for every $y\in\mathcal{N}\setminus V_{2}(\xi_{1},\xi_{2},\dots)$, we have
$\lim_{n\to\infty}\frac{1}{n}\log\left[\big{\|}\mathcal{G}^{\prime}_{\eta,\xi_{n}}(x^{*})\dots\mathcal{G}^{\prime}_{\eta,\xi_{1}}(x^{*})y\big{\|}\right]=\lambda(x^{*}).$
We will call $x^{*}$ linearly stable (or linearly unstable) under stochastic gradient descent if $\lambda(x^{*})<0$ (or $\lambda(x^{*})>0$, respectively). [^8] As mentioned in the introduction, this definition differs from the ones introduced by Wu et al. (2018) and Ma and Ying (2021). See Appendix A for a detailed comparison.
@@ 2.4 Main Result
Recall that, by an argument by Cooper (2021), the Hypothesis (H1) is fulfilled for generic training data, while (H2) holds for the most common choice of initial distribution.
@@
Remark 2.3
Both Theorem A and Theorem B fail to address the case $\mu(x^{*})=0$, respectively $\lambda(x^{*})=0$. Since the support of $X_{\lim}^{GD/SGD}$ is closed by definition and $\mu:\mathcal{M}\to\mathbb{R}$ is continuous, most points with $\mu(x^{*})=0$ should have global minima $x^{\prime}\in\mathcal{M}$ with $\mu(x^{\prime})<0$ nearby and thus be in the support of $X_{\lim}^{GD}$. Making this argument rigorous would require more precise knowledge of the network function $\mathfrak{F}$ and is outside the scope of this paper. Continuity of $\lambda$ is a more subtle issue. In general, $\lambda:\mathcal{M}\to\mathbb{R}$ is only upper semi-continuous [^9] This can readily be seen from the fact that it is defined as an infimum of continuous function.. However, a recent result by Avila et al. (2023) shows that $\lambda$ is continuous at all points $x^{*}$, where the matrices $\mathcal{G}^{\prime}_{\eta,1}(x^{*}),\dots,\mathcal{G}^{\prime}_{\eta,N}(x^{*})$ are all invertible. This is, in particular, the case for all regular $x^{*}\in\mathcal{M}$.
@@ 2.5 Relation to the Neural Tangent Kernel
We point out that, for any $x^{*}\in\mathcal{M}$, the values of $\mu(x^{*})$ and $\lambda(x^{*})$ can be deduced entirely from the neural tangent kernel $K_{x^{*}}:\mathbb{R}^{d}\times\mathbb{R}^{d}\to\mathbb{R}$, first introduced by Jacot et al. (2018), which is defined by
@@
This observation is useful for two reasons. On the one hand it connects the central quantities in the present work to an ongoing line of research of neural networks in the infinite width limit (see Jacot et al. (2018)). On the other hand, it turns out that the expressions (2.18) and (2.19) are nicer to work with than the corresponding expressions in the previous section and will be exclusively used in the proofs of the main theorems.
@@ 2.6 Possible Extensions and Outlook
where $G_{x^{*},[\Xi]}$ denotes the Gram matrix $G_{x^{*}}$ of the neural tangent kernel with all rows whose index is not in $\Xi$ set to 0. In order to derive a version of Theorem B it is necessary to adapt the notion of regular global minima (cf. Definition 2.2) to ensure that the analog of Lemma 3.9 in the proof below still remains valid. We leave this as a problem for future work. The rest of the proof of Theorem B can be extended to mini-batch SGD with only minor changes. Such an extension of Theorem B could help to better understand the impact of the learning rate and the mini-batch size on generalization properties (see e.g. Hoffer et al. (2017); Goyal et al. (2017); Keskar et al. (2017) for numerical studies).
Furthermore it would be interesting to extend our analysis to other versions of stochastic gradient descent such as SGD with momentum (cf. Rumelhart et al. (1986)) or Adam (cf. Kingma and Ba (2014)). The derivations in section 2.3 can be extended to these algorithms by linearizing the iteration steps around fixed points where all moment terms are zero. In the case of Adam, the parameter $\epsilon$ appearing in the numerator of the final update step is usually chosen extremely small (Kingma and Ba (2014) suggest $\epsilon=10^{-8}$). Thus the linearization only approximates the dynamics of the actual algorithm in a vanishingly small neighborhood and it is questionable whether it is still meaningful for the dynamics of the algorithm.
@@
can, even for large $n$, still be negative with small, but positive probability. This effect is captured by a large deviation principle (cf. Arnold and Kliemann (1987)), the rate function of which depends on the moment Lyapunov exponents (see Definition A.2).